### v1

# Minos: Exploiting Cloud Performance Variation with Function-as-a-Service Instance Selection

Abstract.
Serverless Function-as-a-Service (FaaS) is a popular cloud paradigm to quickly and cheaply implement complex applications. Because the function instances cloud providers start to execute user code run on shared infrastructure, their performance can vary. From a user perspective, slower instances not only take longer to complete, but also increase cost due to the pay-per-use model of FaaS services where execution duration is billed with microsecond accuracy. In this paper, we present Minos, a system to take advantage of this performance variation by intentionally terminating instances that are slow. Fast instances are not terminated, so that they can be re-used for subsequent invocations. One use case for this are data processing and machine learning workflows, which often download files as a first step, during which Minos can run a short benchmark. Only if the benchmark passes, the main part of the function is actually executed. Otherwise, the request is re-queued and the instance crashes itself, so that the platform has to assign the request to another (potentially faster) instance. In our experiments, this leads to a speedup of up to 13% in the resource intensive part of a data processing workflow, resulting in up to 4% faster overall performance (and consequently 4% cheaper prices). Longer and complex workflows lead to increased savings, as the pool of fast instances is re-used more often. For platforms exhibiting this behavior, users get better performance and save money by wasting more of the platforms resources.
Serverless, FaaS, workflows, optimization, resource contention
1. Introduction
Figure 1. FaaS platforms start new function instances on shared worker nodes. While users can not influence which worker node will be used, underutilized nodes offer two benefits: faster execution times and lower cost. Using the Minos system, newly started instances check if they are running on a node with low utilization. If not, they terminate themselves by crashing. This leads to a pool of better-performing instances that are re-used for subsequent invocations, leading to a compound performance increase.
Function-as-a-Service (FaaS) offerings have gained increasing popularity in the last years due to their ease of use (Manner, 2023; Bermbach et al., 2017). With FaaS, the cloud platform is responsible for handling most operational tasks, leaving users free to just write the code that should be executed for every incoming event, commonly referred to as functions (Jonas et al., 2019; Manner, 2023). These functions can then be triggered by events in the cloud platform (such as a document upload) or by common transport protocols such as HTTP. In response to incoming events, the cloud platform assigns a function instance (Brooker et al., 2023). If there is no idle already running instance, the platform needs to start a new instance. This process, called a cold start, leads to increased latency for the request as a worker node needs to load the user code and start an isolated environment to run the user code in (Manner et al., 2018). Often, platforms use microVMs (Brooker et al., 2021) or containerization technologies for this (Platfomr, 2020). Users are billed based on the execution time of the invocation, with millisecond accuracy (Manner, 2023). This makes the model cost-efficient for use cases with irregular or infrequent usage patterns, since there is no resource overhead that users need to maintain. As a cloud service hosted on shared infrastructure, functions suffer from performance variations between instances based on the usage patterns of the rest of the platform (Bermbach et al., 2017). For example, as shown in Figure 1, the worker node that the platform selects to start the instance can have different levels of overall utilization. The instance on the worker node with low utilization will benefit from having to share its resources with fewer neighbors. Taking the CPU as example this means that it will have to perform fewer context switches and that it is more likely data can remain cached. These variations also concern most other cloud services. However, FaaS users suffer twice when the variation leads to them having to use slower resources: not only is the execution time longer, but users are also billed more for these slower requests due to the pay-per-use model. In previous work (Schirmer et al., 2023b), we have shown that some FaaS platforms have more than 10% faster performance during the night, when the overall platform is under less load. In this work, we focus on another kind of variation: when starting multiple instances in parallel, some of these instances are going to be faster than other ones. Selectively only running invocations on the fast instances leads to faster and cheaper function executions. To enable this, we present Minos, a system that runs a quick benchmark during every cold start. If the benchmark reveals a particularly slow instance, Minos re-queues the invocation and then terminates the function instance by crashing it. This leads to the platform starting a new instance that might have better performance, or scheduling the invocation on an already running instance. All instances that are already running must have passed the benchmark during their first request (as otherwise, they would have been shut down). This leads to a pool of known-good instances that outperform the average instance, which is especially useful for longer running applications that re-use these instances often. Our evaluation shows that this leads to a state where it is cheaper and faster for users to waste more resources of the platform.
In this paper, we make the following contributions:
  • We propose Minos, a system to take advantage of these variations to select faster instances.
  • We investigate how to choose which functions to keep and which should be terminated.
  • We implement a proof-of-concept prototype for the FaaS platform Google Cloud Functions.
  • We evaluate performance improvements for a data processing use case.
Figure 2. Overview of the process inside a function instance. The first step of the invocation, called prepare, is always executed. If the invocation is a cold start, a benchmark is executed in parallel. Minos then judges whether the result is better than the elysium threshold, in which case nothing happens (and the instance will be re-used in the future). If the instance does not pass, further execution is interrupted, the invocation is re-queued, and the instance is terminated.
2. A System to Exploit Platform Variability
With Minos, [^1] The system is named after the king from Greek mythology, who weights the souls of the dead to decide whether they should ascend to Elysium and enjoy eternal bliss or be thrown into the depths of Tartarus. users put their invocations into a queue, which triggers the platform to execute it. The platform then places it in one of three kinds of execution environments: Either it chooses an already warm instance, which is the ultimate goal of Minos, as the instance has already been identified as having good performance. The platform could also place the invocation on a new instance. In this case, a quick benchmark is performed to measure a metric relevant to the later use case. If the instance passes the benchmark, it will become a known-good instance and be re-used. If not, it is terminated. In order to not lose requests, before terminating, the instance re-queues the invocation that triggered it.
The following sections go into more detail on choosing how much should be terminated (Section 2.1), how to decide on terminating in a distributed manner (Section 2.2), and how the benchmark should work (Section 2.3).
2.1. How much to terminate?
There is a trade-off in how good the performance in the benchmarking step needs to be so that the instance is not terminated. Setting the required performance higher will lead to faster completion times per subsequent request, but it will also lead to many terminated (and subsequently re-queued) invocations, wasting resources. Lower performance requirements lead to shorter completion times and cost in the short term, as requests are not re-queued as often. However, over the long run, the lower average performance of the instances leads to longer completion times per request and thus higher cost. Thus, the optimal termination rate depends on the duration of the workload, the performance variability of the platform, and the relative time (and thus cost) of the benchmark.
$\begin{split}c_{\textsf{total}}=&c_{\textsf{exec}}\left(\sum_{i=0}^{n_{\textsf{term}}}d_{\textsf{term}}+\sum_{i=0}^{n_{\textsf{pass}}}d_{\textsf{pass}}+\sum_{i=0}^{n_{\textsf{reuse}}}d_{\textsf{reuse}}\right)\\ &+\underbrace{c_{\textsf{inv}}\left(n_{\textsf{term}}+n_{\textsf{pass}}+n_{\textsf{reuse}}\right)}_{\approx 0}\end{split}$
Figure 3. The total cost of a workflow comprises the cost per millisecond of execution ($c_{\textsf{exec}}$) of all terminated ($d_{\textsf{term}}$), successful ($d_{\textsf{pass}}$) benchmarks and the re-used instances ($d_{\textsf{reuse}}$), as well as the cost per invocation ($c_{\textsf{inv}}$) per terminated, passed, and re-used function.
As shown in Figure 3, the total cost of a workflow is influenced by the cost per unit of execution and the cost per invocation. In general, the cost per unit of execution dominates the total cost: For the smallest Google Cloud Functions function with 128 MB Memory, the cost per invocation ($c_{\textsf{inv}}$) is roughly equivalent to 50ms of execution time ($c_{\textsf{exec}}$). For the biggest function with 32 GB Memory, it is less than 3 ms. Other platforms follow a similar pricing pattern. Thus, for all longer-running functions, the additional invocations made by Minos are quickly offset by using faster instances.
As a kind of emergency exit, Minos needs to keep track of how often an invocation has been re-queued already. If an invocation has already caused to too many instance terminations compared to the expected rate, this means the platform performance is unusually slow or Minos is unlucky with the instances it gets from the platform right now. In this case, the function is marked as good without performing the benchmark, preventing infinite loops. For example, if the expected termination rate is $40\%$, there is a ca. $1\%$ chance that an invocation fails five times in a row ($0.4^{5}$) and a less than 1‰ chance that the same invocation experiences 8 terminations in a row.
2.2. Autonomous Termination with the Elysium Threshold
It would be optimal to have a centralized component that receives the benchmarking results from all instances to decide what instances are routed requests. While this would allow for complicated scheduling mechanisms, it is not feasible for the kind of highly scalable and elastic workloads that FaaS platforms enable users to run. Instead, a newly started function instance decides locally whether it is good enough based on just one value: what result in the benchmark is necessary to not terminate itself, i.e., so that the function will continue living and working on the next invocations. We call this value the elysium factor. This is stored as part of the function configuration, so that Minos does not require any outside communication during calls.
Calculating the elysium threshold with pre-testing
To calculate the elysium threshold, Minos runs a few benchmarking functions before executing the main workload. This needs to be short enough to not influence the total duration and cost of the overall workflow significantly and may just comprise the first parts of the overall workload running without Minos terminating instances. Based on the performance values, Minos can calculate the elysium threshold to achieve a cost-optimal termination rate and then execute the rest of the workload with Minos terminating instances. As long as the whole workload runs short enough so that longer-term performance variations are not a significant influence (i.e., on the scale of hours), this is a simple and cost-effective approximation of the optimal termination rate.
2.3. What and how long to benchmark
The kind of benchmark that should be performed and the acceptable elysium threshold heavily depends on the use case. Previous research has shown that FaaS workloads are most often constrained by network or CPU resources (Eismann et al., 2021). In our preliminary experiments, CPU performance showed the highest variability and was the bottleneck in most applications.
The benchmark itself should ideally not interfere with the normal execution of the function, i.e., it should run while the resource being benchmarked is otherwise unused. The general approach is shown in Figure 2. During a first step of the execution, when resources need to be set up and downloaded, the benchmark and subsequent judging is run in parallel. For example, in a data processing workflow, the raw data needs to be downloaded and loaded into memory. At the same time, Minos can run its benchmark.
On the other hand, if the function starts with a resource intensive operation, this first operation can be used as benchmark as well. If the kind of operation at the start is representative for the rest of the function, this wastes fewer resources.
Overall, there is not one-size-fits-all solution for what kind of benchmark needs to be executed, and when as well as how long it should run.
3. Evaluation
To demonstrate the feasibility of Minos and evaluate potential performance gains we implement a proof of concept prototype and use it to run a data processing workload.
3.1. Environment Configuration
Use Case
The use case in this evaluation is using linear regression analysis to predict weather. To achieve this, the function first downloads a CSV file containing weather data for a specific location from previous days. Afterwards, the data is analyzed to predict tomorrow’s weather. While downloading the file, Minos can run its CPU benchmark as the download is network-limited. The following linear regression analysis is CPU intensive, benefitting from the instance selection performed in the previous step. To measure the capability of the CPU, we use matrix multiplication as benchmark in Minos (Werner et al., 2018).
Implementation
We implement the prototype to work on Google Cloud Functions. It does not use any features unique to the platform, so that the approach would work on different platforms as well. The function is implemented in Go, a frequently used programming language for serverless applications. The results for the evaluation are read from the function logs after the experiment is finished to rule out influences on execution duration.
Workload
One experiment comprises ten virtual users (VU) sending a request and then waiting one second before sending the next request for 30 minutes. We repeat this experiment every day at the same time for one week. [^2] From 2025-02-03 to 2025-02-09, between 3 and 4 pm UTC As a baseline, we run the same workload at the same time using another function that is exactly the same, except that Minos is completely disabled. All experiments were executed in the europe-west-3 region of Google Cloud Platform with functions configured to have 256 MB RAM (leading to 0.167 vCPU).
Before these runs, we measured the elysium threshold using pre-testing with a short running script where 10 VUs run for one minute (waiting one second between requests). The elysium threshold for the experiment is then set to the 60th percentile of performance we measured (i.e., only the fastest 40% of functions should pass the elysium threshold), which is passed as parameter to the function by the VUs.
For both the baseline and experiment conditions, we measure the time of the function execution, the duration of the benchmark execution, whether it was successful, how often the invocation has been re-tried (retry count), how long the downloading step took (download duration), and how long the linear regression took (analysis duration).
3.2. Results
Linear Regression Duration
Figure 4. The linear regression step was, on average, faster every day the experiment was repeated. The maximum improvement was $>13\%$ on day two, and minimum improvement $4.3\%$ on day three and five. Note that the y-axis is limited to values between 1,000 and 3,000.
The results for the average duration of the computationally heavy linear regression step are shown in Figure 4. Minos lead to better linear regression times on all days the experiment was repeated, with different effect sizes. Over all days, Minos improved the average regression duration by $7.8\%$. This shows that Minos reaches its core goal, reducing the duration of the resource intensive computation.
Successful Requests
Figure 5. Successful requests per day. Minos improved the amount of successful requests all days except one with a maximum of $7.3\%$ on day one, but reduced the amount by $<1\%$ on day five. Note that the y-axis is limited to values between 4,000 and 5,000.
Extending the view from just looking at the resource intensive part of the function, Figure 5 shows how many functions are successfully executed by the ten VUs each day. On the first day, Minos manages to complete 7.3% more requests in the same 30 minutes. This drops in the next days, and at day five, the baseline completed slightly more requests. Over all experiments, Minos completed 2.3% more requests.
Cost
Figure 6. Average Cost per million successful requests per day the experiment ran. On the first and last day, Minos saves more than $3\%$ cost, while it closely tracks the baseline on other days. Note that the y-axis is limited to values between 12 and 14.
We calculate the cost per invocation based on the cost model of Google Cloud Functions. For Minos, we also take the invocations that ultimately were terminated into account and calculate the average cost per actually completed request (successful request). As shown in Figure 6, on day one, Minos saves 3.3% of cost, while also having 3.4% lower latency. This small difference is due to the cost per invocation. On the other days, the cost per successful invocation of Minos is very close to the baseline. This shows that the platform itself is highly variable (as the experiment parameters never changed), and that the performance of Minos depends on the overall cloud platform performance. Over all experiments, Minos was able to reduce the cost per request by 0.9%.
Figure 7. Average cost per million successful requests over all experiments. While the difference is small, Minos has lower cost per successful request for 76% of the experiment duration. Note that the y-axis is limited to values between 10 and 25.
Looking at how the cost develops over the duration of the experiments (cf. Figure 7) shows that, at the beginning, Minos has a higher cost per successful request for the first 200s. This is expected, as Minos terminated many instances, increasing overall cost. Afterwards, the average cost per successful invocation stays quite close, with Minos staying slightly under the baseline for almost the whole duration. After 670 s (11 min 10 s), Minos was cheaper for more than 50 % of time, which increases to 76% at the end of the experiment. This shows that letting Minos run for a longer time increases its benefits.
4. Discussion & Future Work
In the evaluation, Minos showed a small but still considerable improvement over the baseline. Using a data processing application, Minos was able to handle up to 7.3% more request and save up to 3% cost while using more overall resources from the computing platform. This shows that the billing model of Google Cloud Functions, combined with its performance variability, motivates performance and cost oriented users to use more resources of the platform to achieve overall lower cost. With a sufficiently large application, even relatively small savings still lead to considerable cost savings. This poses a problem for the platforms, as their pricing model entices users to save money by wasting platform resources. In general, while the goal of FaaS platforms is to abstract away as much operational complexity as possible, these kinds of applications demonstrate that users still benefit from understanding the platform more deeply.
Cost Models
Future work could focus on researching other cost models that make wasting resources of the cloud provider cost prohibitive. One method could be billing users for the actual CPU performance delivered, not just the theoretical performance, to make less powerful instances financially applicable. This would however require platforms to make it explicit that users can not control the exact pricing and speed of their invocations. This is true today, it is just not represented in the cost model accordingly. Another method would be for platforms to switch to fairer virtualization technologies. For example, the popular FaaS platform AWS Lambda uses microVMs (Brooker et al., 2021), which in our preliminary experiments had a far fairer resource distribution with less overall variability.
Online calculation of the elysium threshold
Our prototype for Minos pre-calculates the elysium threshold. It could also be updated live based on the benchmarking results of currently running instances. After a benchmark has finished, instances can report their results to a centralized component that collects live performance data. It then periodically re-calculates the elysium threshold and updates the functions comparable to our approach in our previous work (Schirmer et al., 2024). While this component is centralized, it is not a single point of failure as its failures would just lead to temporarily suboptimal performance. Depending on the scale of the application, it is not feasible to store all past benchmark results to accurately calculate the performance percentiles. Instead, some statistical properties can be calculated online, i.e., only having to store the most recent new value and a limited amount of information about past results. For example, the mean can be updated online by storing the previous mean and the amount of previous results. Other values, such as the standard deviation (Jain and Chlamtac, 1985) or percentiles (and, 1962) can only be estimated online and not calculated exactly.
Workload Limitations
Note that Minos requires an asynchronous workload, i.e., users need to put their invocations into a queue instead of synchronously waiting for a response. When users require a synchronous response, the instance that receives the request needs to stay alive until the computation is finished to send the response back. If the instance fails the benchmark and should be crashed, it would need to send the call to another function and wait for the response, leading to double billing as the waiting instance still incurs cost (Baldini et al., 2017).
5. Related Work
Overall variability of FaaS functions has, e.g., been shown over longer (Eismann et al., 2022; Ustiugov et al., 2021) and shorter (Schirmer et al., 2023a) time frames, between processor architectures (Lambion et al., 2022), between platforms (Grambow et al., 2024), and for the overall platform behavior (Mahmoudi and Khazaei, 2021; Schirmer et al., 2023b; Ustiugov et al., 2021). Besides compute performance, instance placement can also be used to optimize for fewer carbon emissions (Chadha et al., 2023; Cordingly et al., 2023).
Ginzburg and Freedman present a method for using short-term intra-region FaaS instance variability on AWS Lambda (Ginzburg and Freedman, 2021). They demonstrate temporal, spatial (i.e., using other cloud regions), and instance performance variations in AWS Lambda in 2021. Their evaluation shows that exploiting instance performance variability can lead to performance increases. Compared to Minos, their approach focuses on demonstrating the existence of the effect for CPUs and relies on a centralized scheduler, only working for a limited amount of instances.
Another way of optimizing FaaS applications is speeding up cold starts, e.g., by pre-warming functions before they are called (Bermbach et al., 2020; Bardsley et al., 2018; Carl et al., 2024; Daw et al., 2020). This can be combined with Minos by benchmarking the pre-warmed instances before they are used. Other approaches focus on changing the underlying platform to speed up cold starts, e.g., by changing the virtualization and storage layer (Brooker et al., 2023), or by switching out the scheduling mechanism  (Cvetković et al., 2024).
Another approach is only using FaaS for the parts of the application where its most useful. Other parts can be executed to less expensive compute services (Zhao et al., 2023; Sadeghian et al., 2023).
6. Conclusion
In this paper we have presented Minos, a system to take advantage of FaaS platform variability to save cost and decrease latency. The system benchmarks newly started instances and deliberately terminates them if they do not perform fast enough. This leads to only performant instances being re-used for later invocations, over time leading to a faster pool of instances. Using a prototype built on Google Cloud Functions and a data processing use case, we were able to demonstrate cost savings of up to 3.2% per invocation while handling 7.3% more requests in the same time. Our research shows that cloud platforms need to take precaution to disable these kinds of use cases, as users can waste more platform resources while at the same time paying less overall. Future research can focus on developing fairer virtualization technologies or researching novel cost models.
Acknowledgements.
Partially funded by the Sponsor Bundesministerium für Bildung und Forschung (BMBF, German Federal Ministry of Education and Research) https://www.bmbf.de/bmbf/en in the scope of the Software Campus 3.0 (Technische Universität Berlin) program – Grant #01IS23068.